OpenStudy (anonymous):

12/15= 8/x NEW QUESTION A.9 B. 10 C.12 D.16

OpenStudy (anonymous):

really need help so if u can plz

OpenStudy (otonogold):

You try setting up a proportion and plugging the numbers in?

OpenStudy (anonymous):

what?? like putting a number instead of x?

OpenStudy (otonogold):



OpenStudy (anonymous):

\[\frac{12}{15}=\frac{8}{x}\iff 12x=8\times 15\]

OpenStudy (anonymous):

haha i figured it out it is 10

OpenStudy (otonogold):

Cross multply you will get 120 by 15*8 so then you will divide by 12 getting your answer

OpenStudy (anonymous):

and so \[x=\frac{8\times 15}{12}\]

OpenStudy (otonogold):

good job :)

OpenStudy (anonymous):

thanks for ur guys help :)